OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? A. (1, 7) B. (4, 9) C. (7, 19) D. (8, 21)

OpenStudy (anonymous):

@Abhisar @emfst16 @FelicityJP @LOLAri @LeilaniLane @LinxNeedsHelpPlease @leannaleanna @Write0The0Right0For0Wrong0Doing @E>Nattie12345 canny any of u help

OpenStudy (anonymous):

can

OpenStudy (anonymous):

is there more to the question?

OpenStudy (leilanilane):

Did this question have a chart?

OpenStudy (anonymous):

no thats it

OpenStudy (anonymous):

Try google?

OpenStudy (anonymous):

o wait yes there is 1 sec

OpenStudy (anonymous):

2x-y=-5 x+3y=22

OpenStudy (anonymous):

yea so try plugging in x and y from each and which ever fits the equations is the answer

OpenStudy (leilanilane):

What you're going to do for this problem is find the ordered pair that can solve both equations. All you need to do is substitute the x and y values in both equations.

OpenStudy (leilanilane):

If you want I can help you with one of them.

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

lol i still dont get it (@_o)

OpenStudy (leilanilane):

Okay. So, I'll just start with the first answer, A. (1, 7) x = 1, y = 7 For the first equation: 2x - y = -5 2 * 1 - 7 = - 5 2 - 7 = - 5 -5 = -5 Since both sides (of the equal sign) are... the same, or equal this pair does solve the first equation. For the second equation: x + 3y = 22 1 + 3 * 7 = 22 1 + 21 = 22 22 = 22 Again, both sides are equal. This means this pair is the solution for this equation. I didn't mean to, but I actually found your answer, lol. Its A because its the proper solution for both equations.

OpenStudy (anonymous):

2x1-7=-5 3-7=-5 -5 1+3x7=22 1+21 22

OpenStudy (anonymous):

just plug them in using PEMDAS

OpenStudy (leilanilane):

Do you understand it now? Oh, and if you need more help in the future just let me know. I'm always glad to help if I can. :3

OpenStudy (anonymous):

ok thanks you guys i got it :)

OpenStudy (anonymous):

same here. hope you understand